# Automated training of neural-network interatomic potentials

### v1

@@ Abstract
Neural-network interatomic potentials (NNIPs) have transformed atomistic simulations, by enabling molecular dynamics simulations with near ab initio accuracy at reduced computational costs and improved scalability. Despite these advances, crafting NNIPs remains complex, demanding specialized expertise in both machine learning and electronic-structure calculations. Here, we introduce an automated, open-source, and user-friendly workflow that streamlines the creation of accurate NNIPs. Our approach integrates density-functional theory, data augmentation strategies and classical molecular dynamics to systematically explore the potential energy landscape. Our active-learning strategy leverages on-the-fly calibration of committee disagreement against true errors to ensure reliable uncertainty estimates. We use electronic-structure descriptors and dimensionality reduction to analyze the efficiency of our active learning strategy, which is shown to minimize both false positives and false negatives when deciding what to relabel with ab initio calculations. The method is validated on the fully automated training of a NNIP for a diverse set of carbon allotropes, reaching state-of-the-art accuracy and data efficiency. This platform democratizes NNIP development, empowering users to achieve high-precision simulations with minimal human intervention.
@@
In computational materials science, many key properties—such as phase transitions, diffusion, viscosity, thermal transport—arise from atomic and molecular motion over time. Chemical reactions are typically rare events, requiring long simulations times and enhanced sampling techniques. Because these properties emerge from dynamic processes rather than static configurations, molecular dynamics (MD) simulations are essential for capturing their time-dependent behavior and understanding material performance under different conditions. MD simulations enable the study of crucial phenomena in materials and molecular systems, such as heat conduction, ion transport in batteries, crack propagation in materials, just to name a few, and the behavior of soft matter and biomaterials under different conditions. Fully ab initio molecular dynamics (AIMD) simulations can be highly accurate but computationally expensive, as they require solving the Schrödinger equation every few timesteps to determine atomic forces. This makes AIMD impractical for large systems or long timescales. On the other hand, empirical interatomic potentials, which approximate atomic interactions with predefined functional forms, are computationally efficient but often lack the accuracy needed for quantum mechanical effects. Additionally, these potentials are typically parameterized for specific materials and struggle to generalize across diverse chemical environments, limiting their predictive power for complex or novel systems. For a long time, the Car-Parrinello method (CPMD) [1] offered a compromise between AIMD and empirical potentials by evolving both electronic and ionic degrees of freedom simultaneously with a Lagrangian formalism, reducing the need for explicit electronic structure calculations at every step. While CPMD improves efficiency compared to traditional AIMD, it still remains computationally demanding and is limited by the need for careful tuning of fictitious electron mass parameters, which can affect accuracy and stability.
@@
Over the last years, accuracy and data efficiency of ML interatomic potentials have improved remarkably, often at the price of increased algorithmic complexity. In this context, NNIPs based on equivariance have emerged as particularly promising and several architectures have been proposed, including NequIP [3], Allegro [4] and MACE [5]. More recently, foundation models leveraging large-scale pretraining on diverse chemical datasets, provided a path towards improved transferability, data efficiency, and accuracy across a wide range of materials and molecular systems [6, 7].
As of today, the training of an accurate NNIP remains a complex and time-consuming task. First, high-quality NNIPs require training datasets of the order of thousands of supercell single-point ab initio calculations with hundreds of atoms in the unit cell; millions if foundation models for the entire periodic table are targeted. Most notably, the accuracy and extrapolation capabilities of NNIPs hinge on the careful choice of the training crystal structures, which have to be sufficiently diverse and abundant to avoid overfitting. While foundation models promise transferability across chemical space, fine-tuning them to high accuracy for a given material family still requires curated datasets.
In this work, we introduce AiiDA-TrainsPot, an automated, open-source and user-friendly framework designed to automate the training of NNIPs. Our approach integrates automated workflows for DFT calculations with neural-network training and classical MD to systematically explore the potential energy landscape through random distortions, strain, interfaces, neutral vacancies, trajectories at varying temperatures and pressures. Existing platforms such as DP-GEN [8] or SchNetPack [9] have demonstrated the power of active learning for NNIPs. However, these frameworks are typically tied to specific model families and lack systematic provenance tracking of the training process. In contrast, AiiDA-TrainsPot offers (i) full code-agnostic modularity across quantum engines, ML architectures, and MD codes; (ii) an extensive suite of automated dataset augmentation strategies (defects, slabs, clusters, substitutions); and (iii) a calibrated committee-disagreement scheme that provides quantitative uncertainty estimates even in production runs. The method is validated through the fully automated training of a NNIP for a diverse range of carbon allotropes, achieving state-of-the-art accuracy and data efficiency. Combined with AiiDA’s reproducibility infrastructure, these features uniquely position AiiDA-TrainsPot as both a democratizing tool for domain scientists and a robust platform for future foundation-model development.
I Results
@@
I.1 Automation strategy
AiiDA-TrainsPot is based on a two-stage augmentation process shown in Fig. 1. In a typical use case, users provide a handful of structures—from one up to tens—that are augmented to the order of thousands through structural manipulations. This happens in the first stage, where about a thousand of structures is calculated with ab initio methods (here DFT) and used to train the first generation of NNIPs. The second stage employs MD simulations at different thermodynamic conditions to accurately sample the basins of the potential energy surface (PES), thus exploring regions that are particularly relevant for applications. All MD trajectories are obtained with the NNIPs trained in the previous workflow step. Some structural configurations are sampled from MD trajectories and calculated ab initio to train a second generation of NNIPs. The choice of the structures to label with ab initio results is based on the committee disagreement, i.e., the spread in the predictions of a committee of NNIPs initialized with different seeds and trained on the same data. At each iteration a new generation of potentials is trained and the committee disagreement is calibrated on actual deviations between NNIP predictions and the baseline level of theory, i.e., DFT.
The active learning loop continues until errors on energy, forces and stress tensor are below a user-defined threshold. We emphasize that AiiDA-TrainsPot supports multiple use cases depending on the available input data, which goes beyond the generation of a NNIP from scratch and includes the fine-tuning of foundation models (see Sec. I.2). In the following, we discuss in detail each step of the workflow.
I.1.1 Input structures
@@ I.1.3 Ab Initio Labelling
I.1.4 Training neural-network interatomic potentials
@@
Throughout the active learning iterations, each structure remains in its initially assigned set, ensuring that the test set remains completely independent from the training and validation sets for reliable performance assessment. The training is performed using MACE with default hyperparameters, though these can be fully customized by the user of AiiDA-TrainsPot to suit specific requirements.
I.1.5 Exploration by molecular dynamics
@@ I.1.6 Committee Evaluation
and is not, at least in principle, strictly correlated to the committee disagreement, although empirical evidence suggests they might be linearly related [27] (as we show and discuss more in detail in Sec. I.3 through a validation study on carbon allotropes). After linear regression, the user-defined error tolerance threshold $\epsilon^{thr}_{\alpha}$ (default 1 meV/atom, 100 meV/Å, 1 meV/Å3 for energy, forces and stress tensor respectively) is transformed into an equivalent disagreement threshold $\tau^{thr}_{\alpha}$ that serves as the selection criterion for new structures by $\tau^{thr}_{\alpha}=a_{\alpha}\epsilon^{thr}_{\alpha}$, where coefficients $a_{\alpha}$ are the slopes determined from fitting. The calibration, made at each iteration on the already ab initio labeled structures and compared with $\mathcal{D}_{\alpha}({\mathcal{X}_{i}})$ computed with the last generation potentials, ensures that the uncertainty quantification mechanism remains effective throughout the iterative learning process, as the NNIP committee’s overall accuracy improves with each active learning cycle. Notably, the calibrated committee disagreement $a_{\alpha}\mathcal{D}_{\alpha}({\mathcal{X}_{k}})$ can be used not only in the active learning scheme for the selection of worse predicted structure, but also in production runs, where it can provide an estimate of the uncertainty of the predictions of the NNIPs.
@@ I.2 Code implementation
  • DatasetAugmentationWorkChain for generation of highly-diverse crystal structures
@@
Figure 2: Schematic representation of the TrainsPotWorkChain with its various computational tasks. The workflow begins with the initialization phase, where input structures, parameters, and control flags are set to determine which steps of the workflow will be executed. If dataset augmentation is enabled, the DatasetAugmentationWorkChain generates additional configurations. Next, the workflow enters the active learning loop. Within each iteration, the AbInitioLabellingWorkChain is called to label newly generated configurations by performing automated electronic structure calculations (PwBaseWorkChain). The TrainingWorkChain is then invoked to train NNIPs using MACE (MaceTrainWorkChain). Subsequently, the ExplorationWorkChain executes MD simulations via LAMMPS (LammpsBaseWorkChain) to generate new configurations for further refinement. The EvaluationCalculation assesses the performance of the trained models using calibrated committee disagreement to determine whether additional iterations are required or not.
A key advantage of our implementation strategy and of the use of AiiDA as a workflow engine is the extensive reuse of existing AiiDA plugins, minimizing duplication of software and ensuring robustness through well-tested components. The AbInitioLabellingWorkChain calls PwBaseWorkChain from the AiiDA-Quantum ESPRESSO plugin [28, 29], while the ExplorationWorkChain leverages LammpsBaseWorkChain from the AiiDA-LAMMPS plugin [30]. For specialized functionality not available in existing plugins, we develop custom components such as the DatasetAugmentationWorkChain, which implements various structure manipulation techniques through dedicated calcfunctions based on the Atomic Simulation Environment (ASE) [31]. Similarly, the TrainingWorkChain calls the MaceTrainWorkChain, developed by us, that interfaces with the MACE code for training NNIPs and for handling preprocessing and postprocessing.
@@
Finally, the architecture maintains extensibility through AiiDA’s plugin system, facilitating future integration with additional quantum engines, classical MD codes, or emerging ML frameworks. Individual components like MaceTrainCalculation or EvaluationCalculation can be used as standalone tools outside the high-level WorkChain, making AiiDA-TrainsPot both a comprehensive platform and a flexible toolkit for specialized tasks in interatomic potential development.
@@
I.3 Validation
We showcase AiiDA-TrainsPot with a fully automated training campaign of a NNIP for all carbon allotropes. Carbon represents an ideal test case due to its rich structural diversity, which include 0D fullerenes, 1D nanotubes, 2D graphene, various 3D crystals, including diamond and layered forms such as graphite.
@@ I.3 Validation
The first run, designated as run A or “fast exploration”, is designed to assess the workflow’s capability to explore the PES and study the convergence of the active learning scheme for increasingly large datasets. The run starts by augmenting structures near equilibrium configurations (see Methods Sec. III for more details), that is followed by MD with a wide range of temperatures from 0 up to 5000 K. This also tests the model’s ability to capture extreme conditions such as melting and formation of amorphous phases. As we want to explore convergence for rather large training datasets, which involve up to $10^{4}$ single-point DFT calculations, we employ here the AiiDA-QuantumESPRESSO fast protocol [21] for DFT calculations [28, 29].
The second run, designated as run B or “accuracy and data-efficiency”, focuses, instead, on achieving rather accurate predictions for equilibrium and near-equilibrium conditions with small training datasets (around $2\times 10^{3}$ calculations). The settings of data augmentation and the MD runs are refined for more efficient exploration, including smaller distortions of atomic positions, lower temperatures (up to 1000 K) and larger range of pressures (see Methods Sec. III). As we aim at training an accurate NNIP in an efficient and automated manner, this run employs the default AiiDA-QuantumESPRESSO stringent protocol: not only this ensures more precise DFT reference data, but also reduces noise across the training dataset.
Figure 4: Evolution of model accuracy over active learning. Top panels: RMSE for energies, forces, and stress tensor components are shown across training, validation, and test sets. Bottom panels: Parity plots comparing NNIP predictions (latest iteration of active learning for both runs) to DFT reference values on the final test set. Cold and warm colors indentify the results of run A (“fast exploration”) and run B (“accuracy and data-efficiency”), respectively. Error bars in the top panels represent the standard deviation across the model committee. Insets in the bottom panels show the error distribution histograms.
@@
The fast-exploration campaign runs for 10 iterations of active learning, Fig. 4 reports the evolution of the model performance in cold colors. The top panels track the RMSE for energy, forces, and stress tensor components across training, validation and test sets, as a function of the active learning step and, in parallel, the dataset size that range from the initial 1,177 structures (iteration 1) to 9,537 structures in the final iteration. For energy and stress tensor components, we observe a consistent decrease in prediction errors as the active learning progresses. Interestingly, errors on forces increase from the first to the second iteration, and then decrease monotonically for all the following iterations, suggesting that the active learning strategy first explores novel regions of the PES that require more data to learn. This is supported by the data analysis based on dimensionality reduction discussed later, which shows that early iterations sample an increasing number of distinct structural prototypes. The final RMSE values reach 4.3 meV/atom for energies, 293.0 meV/Å for forces, and 3.7 meV/Å3 for stress tensor components on the test set, comparable to those reported in Ref. [22] for a NNIP trained and tested on all carbon allotropes. The error bars in Fig. 4 represent the standard deviation across the model committee, which also decrease with iterations as the model becomes more consistent and robust. The bottom panels display parity plots comparing DFT reference values against NNIP predictions for the test set after the final iteration. The tight clustering of points along the diagonal line, together with the error distribution histograms (insets), demonstrates excellent agreement between the ML predictions and DFT calculations across all evaluated properties.
@@
Beyond understanding the exploration strategy through SOREP analysis, we evaluate the effectiveness of using committee disagreement as an uncertainty metric for steering the growth of the training dataset. Figure 5b demonstrates the correlation between committee disagreement $\mathcal{D}_{\alpha}(\mathcal{X}_{k})$ and true prediction errors $\epsilon_{\alpha}(\mathcal{X}_{k})$ at the final active learning iteration. The analysis reveals an approximately linear relationship between these quantities, confirming that committee disagreement can serve as a reliable proxy for accuracy [27], although the proportionality constant is far from being unity. As anticipated in Sec. I.1.6, we introduce a calibration factor $a_{\alpha}$ determined by linear regression, which transforms user-defined error tolerances $\epsilon^{thr}_{\alpha}$ into equivalent disagreement thresholds $\tau^{thr}_{\alpha}$; the calibrated committee disagreement is then used to select the structures to calculate from first principles. This calibration procedure ensures that the committee disagreement threshold appropriately reflects the true prediction errors, addressing cases where the uncalibrated disagreement metric might either overestimate (as observed for forces in Fig. 5b) or underestimate (as seen for energies and stress in our validation test) the actual deviations from DFT.
@@
While the first validation study demonstrates the effectiveness of our active learning strategy, we observe diminishing returns in RMSE improvement as the number of iterations and the size of the dataset increase, due to the exploration of larger and larger regions of the PES. This suggests that strategic optimization of initial structures and data augmentation parameters can enhance the NNIP accuracy for the target application, while performing very few iterations of active learning. Therefore, for the second validation—focused on accuracy and data efficiency—we target the description of polymorphs near equilibrium with a refined data augmentation strategy combined with constrained temperature ranges in MD simulations: strain range is increased with respect to the previous run, while atomic rattling is reduced and MD temperatures range up to 1000 K (see Methods III for details). This approach delivers high-quality potentials in just two active learning iterations and about 1800 training configurations, hence making more affordable and sustainable the use of the stringent protocol for DFT calculations, which is computationally more expensive.
@@
The left panel of Fig. 7 shows the excellent agreement between NNIP predictions and DFT references in around equilibrium; discrepancies become more noticeable at extreme bond compressions or expansions, which correspond to configurations that are underrepresented in the training data. Notably, the EOS for the dimer is reasonably accurate, even if no dimer configurations were included in the initial training set. The right-hand panel compares the EOS for graphite, graphene, and diamond, focusing on their relative energetic ordering. This inset is especially informative because graphite and diamond exhibit nearly degenerate formation energies in DFT. The NNIP correctly reproduces the energy hierarchy, demonstrating its ability to capture subtle thermodynamic trends. Similar benchmarking practices have been applied in the development of the ACE family of potentials [22].

### v2

# AiiDA-TrainsPot: Towards automated training of neural-network interatomic potentials

@@ Abstract
Crafting neural-network interatomic potentials (NNIPs) remains a complex task, demanding specialized expertise in both machine learning and electronic-structure calculations. Here, we introduce AiiDA-TrainsPot, an automated, open-source, and user-friendly workflow that streamlines the creation of accurate NNIPs by orchestrating density-functional-theory calculations, data augmentation strategies, and classical molecular dynamics. Our active-learning strategy leverages on-the-fly calibration of committee disagreement against ab initio reference errors to ensure reliable uncertainty estimates. We use electronic-structure descriptors and dimensionality reduction to analyze the efficiency of this calibrated criterion, and show that it minimizes both false positives and false negatives when deciding what to compute from first principles. AiiDA-TrainsPot has a modular design that supports multiple NNIP backends, enabling both the training of NNIPs from scratch and the fine-tuning of foundation models. We demonstrate its capabilities through automated training campaigns targeting pristine and defective carbon allotropes, including amorphous carbon, as well as structural phase transitions in monolayer $\mathrm{W_{x}Mo_{1-x}Te_{2}}$ alloys.
I Introduction
@@
In computational materials science, many key properties—such as phase transitions, diffusion, viscosity, thermal transport—arise from atomic and molecular motion over time. Chemical reactions are typically rare events, requiring long simulations times and enhanced sampling techniques. Because these properties emerge from dynamic processes rather than static configurations, molecular dynamics (MD) simulations are essential for capturing their time-dependent behavior and understanding material performance under different conditions. MD simulations enable the study of crucial phenomena in materials and molecular systems, such as heat conduction, ion transport in batteries, crack propagation in materials, just to name a few, and the behavior of soft matter and biomaterials under different conditions. Fully ab initio molecular dynamics (AIMD) simulations can be highly accurate but computationally expensive, as they require solving the Schrödinger equation every timestep to determine atomic forces. This makes AIMD impractical for large systems or long timescales. On the other hand, empirical interatomic potentials, which approximate atomic interactions with predefined functional forms, are computationally efficient but often lack the accuracy needed for quantum mechanical effects. Additionally, these potentials are typically parameterized for specific materials and struggle to generalize across diverse chemical environments, limiting their predictive power for complex or novel systems. For a long time, the Car-Parrinello method (CPMD) [17] offered a compromise between AIMD and empirical potentials by evolving both electronic and ionic degrees of freedom simultaneously with a Lagrangian formalism, reducing the need for explicit electronic structure calculations at every step. While CPMD improves efficiency compared to traditional AIMD, it still remains computationally demanding and is limited by the need for careful tuning of fictitious electron mass parameters, which can affect accuracy and stability.
@@
Over the last years, accuracy and data efficiency of ML interatomic potentials have improved remarkably, often at the price of increased algorithmic complexity. In this context, NNIPs based on equivariance have emerged as particularly promising and several architectures have been proposed, including NequIP [8], Allegro [44], MACE [7] and Point Edge Transformer (PET) [50]. More recently, foundation models leveraging large-scale pretraining on diverse chemical datasets, provided a path towards improved transferability, data efficiency, and accuracy across a wide range of materials and molecular systems [37, 42].
As of today, the training of an accurate NNIP remains a complex and time-consuming task. First, high-quality NNIPs require training datasets of the order of thousands of supercell single-point ab initio calculations with hundreds of atoms in the unit cell; millions if foundation models for the entire periodic table are targeted. Most notably, the accuracy and extrapolation capabilities of NNIPs hinge on the careful choice of the training structures, which have to be sufficiently diverse and abundant to avoid overfitting. While foundation models promise transferability across chemical space, fine-tuning them to high accuracy for a given material family still requires curated datasets.
In this work, we introduce AiiDA-TrainsPot, an automated, open-source and user-friendly framework designed to automate the training of NNIPs. Our approach integrates automated workflows for DFT calculations with neural-network training and classical MD to systematically explore the potential energy landscape through random distortions, strain, interfaces, neutral vacancies, trajectories at varying temperatures and pressures. Existing platforms such as DP-GEN [74] or SchNetPack [59] have demonstrated the power of active learning for NNIPs. However, most existing frameworks are tied to specific NNIP architectures or training backends, offer only limited error handling for first-principles simulations, and do not provide systematic provenance tracking of the training process. In contrast, AiiDA-TrainsPot offers (i) full code-agnostic modularity across quantum engines, ML architectures, and MD codes; (ii) an extensive suite of automated dataset augmentation strategies (defects, slabs, clusters, substitutions, alloys); and (iii) a calibrated committee-disagreement scheme that provides quantitative uncertainty estimates even in production runs. The method is validated through the fully automated training and fine-tuning of NNIPs for a diverse range of carbon allotropes, including amorphous carbon, as well as for monolayer $\mathrm{W_{x}Mo_{1-x}Te_{2}}$ alloys, achieving state-of-the-art accuracy and data efficiency. Combined with AiiDA’s reproducibility infrastructure, these features uniquely position AiiDA-TrainsPot as both a democratizing tool for domain scientists and a robust platform for future foundation-model development.
II Results and discussion
@@
II.1 Automation strategy
AiiDA-TrainsPot is based on a two-stage augmentation process shown in Fig. 1. In a typical use case, users provide a handful of structures—from one up to tens—that are augmented to the order of thousands through structural manipulations. This happens in the first stage, where about a thousand of structures are calculated with ab initio methods (here DFT) and used to train the first generation of NNIPs. The second stage employs MD simulations at different thermodynamic conditions to accurately sample the basins of the potential energy surface (PES), thus exploring regions that are particularly relevant for applications. All MD trajectories are obtained with the NNIPs trained in the previous workflow step. Some structural configurations are sampled from MD trajectories and calculated ab initio to train a second generation of NNIPs. The choice of the structures to label with ab initio results is based on the committee disagreement, i.e., the spread in the predictions of a committee of NNIPs initialized with different seeds and trained on the same data. At each iteration a new generation of potentials is trained and the committee disagreement is calibrated on actual deviations between NNIP predictions and the baseline level of theory, i.e., DFT.
The active learning loop continues until errors on energy, forces and stress tensor are below a user-defined threshold. We emphasize that AiiDA-TrainsPot supports multiple use cases depending on the available input data, which goes beyond the generation of a NNIP from scratch and includes the fine-tuning of foundation models (see Sec. II.2). In the following, we discuss in detail each step of the workflow.
II.1.1 Input structures
@@ II.1.2 Dataset augmentation
  • Alloys: Alloy configurations are generated by randomly mixing atomic species (specified via alloy_species), while optionally keeping some species fixed (specified via fixed_species). If no target compositions are specified, the workflow samples random alloy configurations with concentrations spanning the full range from 0 to 1.
@@ II.1.3 Ab Initio Labelling
II.1.4 Training neural-network interatomic potentials
@@
Throughout the active learning iterations, each structure remains in its initially assigned set, ensuring that the test set remains completely independent from the training and validation sets for reliable performance assessment. The training is performed using either MACE or Metatrain [13] with default hyperparameters, though these can be fully customized by the user of AiiDA-TrainsPot to suit specific requirements.
II.1.5 Exploration by molecular dynamics
@@ II.1.6 Committee Evaluation
and is not, at least in principle, strictly correlated to the committee disagreement, although empirical evidence suggests they might be linearly related [35] (as we show and discuss more in detail in Sec. II.3 through a validation study on carbon allotropes). After linear regression, the user-defined error tolerance threshold $\epsilon^{thr}_{\alpha}$ (default 1 meV/atom, 100 meV/Å, 1 meV/Å3 for energy, forces and stress tensor respectively) is transformed into an equivalent disagreement threshold $\tau^{thr}_{\alpha}$ that serves as the selection criterion for new structures by $\tau^{thr}_{\alpha}=a_{\alpha}\epsilon^{thr}_{\alpha}$, where coefficients $a_{\alpha}$ are the slopes determined from fitting. The calibration, made at each iteration on the already ab initio labeled structures and compared with $\mathcal{D}_{\alpha}({\mathcal{X}_{i}})$ computed with the last generation potentials, ensures that the uncertainty quantification mechanism remains effective throughout the iterative learning process, as the NNIP committee’s overall accuracy improves with each active learning cycle. Notably, the calibrated committee disagreement $a_{\alpha}\mathcal{D}_{\alpha}({\mathcal{X}_{k}})$ can be used not only in the active learning scheme for the selection of worse predicted structure, but also in production runs, where it can provide an estimate of the uncertainty of the predictions of the NNIPs.
@@ II.2 Code implementation
  • DatasetAugmentationWorkChain for generation of highly-diverse training structures
@@
Figure 2: Schematic representation of the TrainsPotWorkChain with its various computational tasks. The workflow begins with the initialization phase, where input structures, parameters, and control flags are set to determine which steps of the workflow will be executed. If dataset augmentation is enabled, the DatasetAugmentationWorkChain generates additional configurations. Next, the workflow enters the active learning loop. Within each iteration, the AbInitioLabellingWorkChain is called to label newly generated configurations by performing automated electronic structure calculations (PwBaseWorkChain). The TrainingWorkChain is then invoked to train interatomic potentials using MACE (MaceTrainWorkChain) or Metatrain (MetaTrainWorkChain). Subsequently, the ExplorationWorkChain executes MD simulations via LAMMPS (LammpsBaseWorkChain) to generate new configurations for further refinement. The EvaluationCalculation assesses the performance of the trained models using calibrated committee disagreement to determine whether additional iterations are required or not.
A key advantage of our implementation strategy and of the use of AiiDA as a workflow engine is the extensive reuse of existing AiiDA plugins, minimizing duplication of software and ensuring robustness through well-tested components. The AbInitioLabellingWorkChain calls PwBaseWorkChain from the AiiDA-Quantum ESPRESSO plugin [33, 68], while the ExplorationWorkChain leverages LammpsBaseWorkChain from the AiiDA-LAMMPS plugin [1]. For specialized functionality not available in existing plugins, we develop custom components such as the DatasetAugmentationWorkChain, which implements various structure manipulation techniques through dedicated calcfunctions based on the Atomic Simulation Environment (ASE) [39]. Similarly, the TrainingWorkChain interfaces with NN and ML training codes and can be configured by the user to choose the training instrument, either MaceTrainWorkChain for MACE or MetatrainWorkChain for Metatrain, handling all preprocessing, training, and postprocessing steps in a fully automated manner. We note that Metatrain currently implements both the PET architecture and other ML architectures not all based on neural-networks, such as Sparse Gaussian Approximation Potentials (GAP) [5] and Behler-Parrinello neural networks with SOAP features (SOAP BPNN) [10].
@@
Finally, the architecture maintains extensibility through AiiDA’s plugin system, facilitating future integration with additional quantum engines, classical MD codes, or emerging ML frameworks. Individual components like MaceTrainCalculation, MetaTrainCalculation or EvaluationCalculation can be used as standalone tools outside the high-level WorkChain, making AiiDA-TrainsPot both a comprehensive platform and a flexible toolkit for specialized tasks in interatomic potential development.
@@
II.3 Validation
To validate our workflow and demonstrate its capabilities, we consider two representative case studies. First, we carry out a fully automated training campaign of a NNIP for a comprehensive set of carbon allotropes, showcasing the ability of the workflow to handle materials with widely different bonding topologies and dimensionalities. Second, we investigate 2D transition-metal dichalcogenides, focusing on monolayer $\mathrm{W_{x}Mo_{1-x}Te_{2}}$ alloys, to illustrate the flexibility of the approach for multicomponent systems and its accuracy in capturing structural phase transitions as a function of composition.
Carbon allotropes
As a first validation study, we showcase AiiDA-TrainsPot with an active learning training of a NNIP for all carbon allotropes. Carbon represents an ideal test case due to its rich structural diversity, which includes 0D fullerenes, 1D nanotubes, 2D graphene, various 3D crystals, including diamond and layered forms such as graphite.
@@ II.3 Validation
The first run, designated as run A or “fast exploration”, is designed to assess the workflow’s capability to explore the PES and study the convergence of the active learning scheme for increasingly large datasets. The run starts by augmenting structures near equilibrium configurations (see Methods Sec. IV for more details), that is followed by MD with a wide range of temperatures from 0 up to 5000 K. This also tests the model’s ability to capture extreme conditions such as melting and formation of amorphous phases. As we want to explore convergence for rather large training datasets, which involve up to $10^{4}$ single-point DFT calculations, we employ here the AiiDA-QuantumESPRESSO fast protocol [45] for DFT calculations [33, 68].
The second run, designated as run B or “accuracy and data-efficiency”, focuses, instead, on achieving rather accurate predictions for equilibrium and near-equilibrium conditions with small training datasets (around $2\times 10^{3}$ calculations). The settings of data augmentation and the MD runs are refined for more efficient exploration, including smaller distortions of atomic positions, lower temperatures (up to 1000 K) and larger range of pressures (see Methods Sec. IV). As we aim at training an accurate NNIP in an efficient and automated manner, this run employs the default AiiDA-QuantumESPRESSO stringent protocol: not only this ensures more precise DFT reference data, but also reduces noise across the training dataset.
Figure 4: Evolution of model accuracy over active learning. Top panels: RMSE for energies, forces, and stress tensor components are shown across training, validation, and test sets. Bottom panels: Parity plots comparing NNIP predictions (latest iteration of active learning for both runs) to DFT reference values on the final test set. Cold and warm colors identify the results of run A (“fast exploration”) and run B (“accuracy and data-efficiency”), respectively. Error bars in the top panels represent the standard deviation across the model committee. Insets in the bottom panels show the error distribution histograms.
@@
The fast-exploration campaign runs for 10 iterations of active learning, Fig. 4 reports the evolution of the model performance in cold colors. The top panels track the RMSE for energy, forces, and stress tensor components across training, validation and test sets, as a function of the active learning step and, in parallel, the dataset size that range from the initial 1,177 structures (iteration 1) to 9,537 structures in the final iteration. For energy and stress tensor components, we observe a consistent decrease in prediction errors as the active learning progresses. Interestingly, errors on forces increase from the first to the second iteration, and then decrease monotonically for all the following iterations, suggesting that the active learning strategy first explores novel regions of the PES that require more data to learn. This is supported by the data analysis based on dimensionality reduction discussed later, which shows that early iterations sample an increasing number of distinct structural prototypes. The final RMSE values reach 4.3 meV/atom for energies, 293.0 meV/Å for forces, and 3.7 meV/Å3 for stress tensor components on the test set, comparable to those reported in Ref. [53] for a ML potential trained and tested on all carbon allotropes. The error bars in Fig. 4 represent the standard deviation across the model committee, which also decrease with iterations as the model becomes more consistent and robust. The bottom panels display parity plots comparing DFT reference values against NNIP predictions for the test set after the final iteration. The tight clustering of points along the diagonal line, together with the error distribution histograms (insets), demonstrates excellent agreement between the ML predictions and DFT calculations across all evaluated properties.
@@
Beyond understanding the exploration strategy through SOREP analysis, we evaluate the effectiveness of using committee disagreement as an uncertainty metric for steering the growth of the training dataset. Figure 5b demonstrates the correlation between committee disagreement $\mathcal{D}_{\alpha}(\mathcal{X}_{k})$ and true prediction errors $\epsilon_{\alpha}(\mathcal{X}_{k})$ at the final active learning iteration. The analysis reveals an approximately linear relationship between these quantities, confirming that committee disagreement can serve as a reliable proxy for accuracy [35], although the proportionality constant is far from being unity. As anticipated in Sec. II.1.6, we introduce a calibration factor $a_{\alpha}$ determined by linear regression, which transforms user-defined error tolerances $\epsilon^{thr}_{\alpha}$ into equivalent disagreement thresholds $\tau^{thr}_{\alpha}$; the calibrated committee disagreement is then used to select the structures to calculate from first principles. This calibration procedure ensures that the committee disagreement threshold appropriately reflects the true prediction errors, addressing cases where the uncalibrated disagreement metric might either overestimate (as observed for forces in Fig. 5b) or underestimate (as seen for energies and stress in our validation test) the actual deviations from DFT.
@@
While the first validation study demonstrates the effectiveness of our active learning strategy, we observe diminishing returns in RMSE improvement as the number of iterations and the size of the dataset increase, due to the exploration of larger and larger regions of the PES. This suggests that strategic optimization of initial structures and data augmentation parameters can enhance the NNIP accuracy for the target application, while performing very few iterations of active learning. Therefore, for the second validation—focused on accuracy and data efficiency—we target the description of polymorphs near equilibrium with a refined data augmentation strategy combined with constrained temperature ranges in MD simulations: strain range is increased with respect to the previous run, while atomic rattling is reduced and MD temperatures range up to 1000 K (see Methods IV for details). This approach delivers high-quality potentials in just two active learning iterations and about 1800 training configurations, hence making more affordable and sustainable the use of the stringent protocol for DFT calculations, which is computationally more expensive.
@@
The left panel of Fig. 7 shows the excellent agreement between NNIP predictions and DFT references in around equilibrium; discrepancies become more noticeable at extreme bond compressions or expansions, which correspond to configurations that are underrepresented in the training data. Notably, the EOS for the dimer is reasonably accurate, even if no dimer configurations were included in the initial training set. The right-hand panel compares the EOS for graphite, graphene, and diamond, focusing on their relative energetic ordering. This inset is especially informative because graphite and diamond exhibit nearly degenerate formation energies in DFT. The NNIP correctly reproduces the energy hierarchy, demonstrating its ability to capture subtle thermodynamic trends. Similar benchmarking practices have been applied in the development of the ACE interatomic potentials [53].